OpenStudy (anonymous):

Select thepoint that lies on the graph of y < 2x -5. Choose one answer. a. (-1, 7) b. (0, 3) c. (1, 3) d. (3, -6)

OpenStudy (anonymous):

please help! :)

OpenStudy (anonymous):

y = -6 2x-5 = 6 - 5 = 1 so y < 2x-5 answer is d

OpenStudy (anonymous):

did you follow that ok?

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

thats ok - if you are happy with the answer plz click on good answer box to close the question